# BehaviorGPT: Smart Agent Simulation for Autonomous Driving with Next-Patch Prediction

### v1

@@ 3.5 Next-Patch Prediction Head
Given the semantically rich and interaction-aware patch features output by the Transformer decoder, we develop a next-patch prediction head to model the multimodal distribution of agent trajectories. We employ an autoregressive masking mechanism during inference to sample agents’ trajectories patch by patch. Specifically, based on the attention output of a patch, the prediction head will estimate the distribution of the next patch. After sampling the next trajectory patch from the distribution, we concatenate the predicted patch with the original sequence and continue to reason about the subsequent patches. An example with a patch size of is shown in Figure 4.
The following describes the process of next-patch prediction regarding the -th patch of the -th agent. Given the attention output , we intend to estimate the parameters of the next patch’s mixture model, which we pre-define with modes. To begin with, we use an MLP to transform into , i.e., the mixing coefficient of the modes. On the other hand, we employ a GRU-based autoregressive RNN [10] to unroll the states within the next patch step by step, which respects the sequential nature of behavior simulation. Specifically, the hidden state of the RNN is initialized with and is updated by the newly predicted states step by step during the rollout. At each step of the RNN, we use an MLP to estimate the location and scale parameters of the next agent state’s distribution based on the hidden state. The whole process is formulated as follows:
where and are locations and scales. We consider the distribution of the next agent state as a multivariate marginal distribution, parameterizing the position and velocity components as Laplace distributions and the yaw angle as a Von Mises distribution.
@@ 3.6 Training Objective
To train BehaviorGPT, we employ the negative log-likelihood loss on the factorized as formulated previously:
Note that each ground-truth trajectory patch is transformed into the viewpoint of its previous patch. During training, we utilize teacher forcing to parallelize the modeling of next-patch prediction and lower the learning difficulty, but we do not use the ground-truth agent states when updating the RNN’s hidden states, intending to train the model to recover from its mistakes made in next-state prediction.
@@ 4 Experiments
We use the Waymo Open Motion Dataset (WOMD) [16] to evaluate the effectiveness of our BehaviorGPT. We first introduce the dataset and the evaluation metrics, then present the rollout results obtained by BehaviorGPT on the Waymo Sim Agents Benchmark. Finally, we conduct ablation study to further compare and analyze the performance of BehaviorGPT under various settings.
@@ 4.1 Dataset and Metrics
Our experiments are conducted on the WOMD v1.2.1. The dataset comprises 486,995/44,097/44,920 training/validation/testing scenarios. Each scenario includes -step historical observations and -step future observations, totaling seconds. We simulate up to agents and generate simulation steps ( seconds) per agent at -second intervals in an autoregressive and reactive manner for evaluation. Each agent requires simulations, each including x/y/z centroid coordinates and a heading value. We use NVIDIA RTX 4090 GPUs to train models, where using training data takes hours and using training data takes days.
@@
We use various metrics to evaluate our method. The minADE measures the minimum average displacement error over multiple simulated trajectories, assessing trajectory accuracy. REALISM is the meta metric that expects the simulations to match real-world observations. LINEAR ACCEL and LINEAR SPEED evaluate the realism regarding acceleration and speed. ANG ACCEL and ANG SPEED measure angular acceleration and speed used in maneuvering. DIST TO OBJ assesses the distances to objects, while COLLISION and TTC evaluate collision frequency and time to collision. Finally, DIST TO ROAD and OFFROAD measure map compliance.
4.2 Results on Waymo Sim Agents Benchmark
We evaluated the performance of our model using the Waymo Sim Agents Benchmark [30] and reported the results in Table 1. Although there were minor changes in the official calculation of metrics such as LINEAR ACCEL and COLLISION in 2024, the minADE metric remains stable and consistent. Thus, we consider minADE to be the primary metric for comparison. Our performance surpasses the state-of-the-art Trajeglish [33] across multiple metrics while also outperforming several well-established motion prediction models, including Wayformer [31], MultiPath++ [44], and MTR [40]. Notably, BehaviorGPT achieves the lowest minADE, indicating superior trajectory accuracy. Its high realism score underscores the model’s ability to match the real-world distribution while also maintaining competitive performance in speed, acceleration, and collision metrics. These strengths indicate that BehaviorGPT is a robust and reliable choice for traffic agent simulation, showcasing its advanced capabilities in simulating agents’ behavior in dynamic environments. As of the time we submitted the paper, the Waymo Open Sim Agents Challenge 2024 had not ended. We will update the complete results on the leaderboard in the revised version of our paper.
@@ 4.2 Results on Waymo Sim Agents Benchmark
Model | minADE () | REALISM | LINEAR ACCEL | LINEAR SPEED | ANG ACCEL | ANG SPEED | DIST TO OBJ | COLLISION | TTC | DIST TO ROAD | OFFROAD |
Random Agent [30] | 50.706 | 0.155 | 0.044 | 0.002 | 0.012 | 0.074 | 0.000 | 0.000 | 0.734 | 0.178 | 0.287 |
Wayformer [31] | 2.498 | 0.575 | 0.098 | 0.331 | 0.406 | 0.413 | 0.297 | 0.870 | 0.782 | 0.592 | 0.866 |
CAD [9] | 2.308 | 0.531 | 0.253 | 0.349 | 0.310 | 0.432 | 0.332 | 0.568 | 0.789 | 0.637 | 0.834 |
MultiPath++ [44] | 2.049 | 0.533 | 0.230 | 0.434 | 0.452 | 0.515 | 0.345 | 0.567 | 0.812 | 0.639 | 0.682 |
MVTA [47] | 1.866 | 0.636 | 0.220 | 0.439 | 0.480 | 0.533 | 0.374 | 0.875 | 0.829 | 0.654 | 0.893 |
MTR+++ [40] | 1.679 | 0.608 | 0.107 | 0.414 | 0.436 | 0.484 | 0.347 | 0.861 | 0.797 | 0.654 | 0.895 |
MVTE [47] | 1.674 | 0.645 | 0.222 | 0.445 | 0.481 | 0.535 | 0.383 | 0.893 | 0.832 | 0.664 | 0.908 |
Trajeglish [33] | 1.571 | 0.645 | 0.192 | 0.453 | 0.538 | 0.485 | 0.386 | 0.922 | 0.836 | 0.659 | 0.886 |
Waymo Baseline (2024) [30] | 7.514 | 0.398 | 0.116 | 0.043 | 0.439 | 0.252 | 0.215 | 0.390 | 0.755 | 0.480 | 0.442 |
MVTE (2024) [47] | 1.676 | 0.730 | 0.353 | 0.359 | 0.599 | 0.497 | 0.374 | 0.904 | 0.830 | 0.665 | 0.907 |
BehaviorGPT (Ours) | 1.540 | 0.741 | 0.315 | 0.355 | 0.530 | 0.469 | 0.377 | 0.952 | 0.828 | 0.666 | 0.930 |
Besides the benchmarking results, we further compare the parameter count of BehaviorGPT against several baselines. Table 2 demonstrates that BehaviorGPT, with only 3 million parameters, achieves a notable reduction in parameter count compared with state-of-the-art models like MTR++ and MVTA. When compared against Trajeglish, BehaviorGPT achieves an approximately 91.6% reduction in parameter count. Despite its minimal parameter count, BehaviorGPT outperforms these more complex models, as shown in Table 1. These results demonstrate the parameter efficiency of decoder-only architecture and also indicate that our approach can utilize the training samples more efficiently.
4.3 Impact of Hyperparameters
In BehaviorGPT, four model hyperparameters are expected to impact model performance: the patch size, the number of spatial neighbors, the number of modes, and the probability threshold for trajectory mode sampling. We trained and evaluated BehaviorGPT using different hyperparameters to investigate their impact on model performance in the following.
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (patch size=1) | 2.3752 | 0.6783 | 0.2088 | 0.4021 | 0.3201 | 0.8432 |
BehaviorGPT (patch size=5) | 1.5598 | 0.7272 | 0.3217 | 0.4622 | 0.3768 | 0.9077 |
BehaviorGPT (patch size=10) | 1.5203 | 0.7334 | 0.3022 | 0.4733 | 0.3796 | 0.9131 |
@@ 4.3 Impact of Hyperparameters
Table 3 presents the results of BehaviorGPT with varying patch sizes. By comparing the most representative metrics, it is evident that using a patch size of , meaning training and predicting with steps ( seconds), significantly outperforms the model without patching. Additionally, increasing the patch size to further enhances the performance, though the improvement is less substantial. These results demonstrate the benefits of incorporating the NP3 into agent simulation.
The next hyperparameter to be investigated is the number of spatial neighbors considered in Transformers. Table 4 shows that increasing the neighbor number from to can improve the performance on all metrics, which shows that our model can easily scale up with more computation.
@@
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (k=16) | 1.6463 | 0.7207 | 0.2659 | 0.4510 | 0.6540 | 0.8926 |
BehaviorGPT (k=32) | 1.5930 | 0.7283 | 0.2691 | 0.4556 | 0.6550 | 0.9137 |
Lastly, the number of modes refers to the number of predicted futures in multi-modal distributions, where a larger number of modes may enhance the trajectory diversity. We also adopt top- sampling [21] to select the smallest set of most probable tokens whose cumulative probability exceeds , where a larger encourages more diverse outputs. Table 5 reveals that for the same top- value, increasing the number of modes () generally improves the realism and offroad metrics but has varied effects on other metrics. For the same , increasing achieves better realism and linear acceleration but slightly decreases minADE and other metrics.
@@
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (=8, p=0.95) | 1.6132 | 0.7290 | 0.2630 | 0.4564 | 0.6520 | 0.9174 |
BehaviorGPT (=16, p=0.8) | 1.6278 | 0.7274 | 0.2652 | 0.4557 | 0.6547 | 0.9148 |
BehaviorGPT (=16, p=0.9) | 1.5671 | 0.7298 | 0.2791 | 0.4631 | 0.6606 | 0.9147 |
BehaviorGPT (=16, p=0.95) | 1.5479 | 0.7310 | 0.2883 | 0.4686 | 0.6625 | 0.9131 |
BehaviorGPT (=16, p=1.0) | 1.5276 | 0.7335 | 0.2959 | 0.4744 | 0.6630 | 0.9142 |
@@ 5 Conclusion
In this work, we introduced BehaviorGPT, a decoder-only autoregressive architecture designed to enhance smart agent simulation for autonomous driving. By applying homogeneous Transformer blocks to entire trajectory snippets and utilizing relative spacetime representations, BehaviorGPT simplifies the modeling process and maximizes data utilization. To enable high-level understanding and long-range interaction reasoning in space and time, we developed the Next-Patch Prediction Paradigm, which requires the model to predict patches of trajectories instead of short segments. Our results on the Waymo Open Sim Agents Challenge demonstrate that BehaviorGPT, with just 3 million parameters, achieves superior performance, highlighting its potential to improve the realism and scalability of autonomous vehicle simulations with significantly fewer parameters.

### v2

@@ 3.5 Next-Patch Prediction Head
Given the interaction-aware patch features output by the Transformer decoder, we develop a next-patch prediction head to model the marginal multimodal distribution of agent trajectories, which estimates the distributional parameters of each patch’s successor.
The following describes the process of next-patch prediction regarding the -th patch of the -th agent. Based on the attention output , we intend to estimate the parameters of the next patch’s mixture model pre-defined with modes. First, we use an MLP to transform into , the mixing coefficient of the modes. In each mode, the conditional distribution of the next agent state, as depicted in Eq. (5), is considered a multivariate marginal distribution that parameterizes the position and velocity components as Laplace distributions and the yaw angle as a von Mises distribution. Based on this formulation, we employ a GRU-based autoregressive RNN [9] to unroll the states within the next patch step by step, with each step being conditioned on the previously predicted states. Specifically, The hidden state of the RNN is initialized with at for . At each step of the rollout, we use an MLP to estimate the location and scale parameters of the next agent state’s position and velocity based on the hidden state. On the other hand, the MLP also estimates the location and concentration parameters of the next yaw angle. The location parameters of the newly predicted state, including the 3D positions, the 2D velocities, and the yaw angle, are used to update the RNN’s hidden state directly without relying on the predicted scale/concentration parameters for sampling. The whole process is summarized as follows:
where , , and are the location, scale, and concentration parameters in the -th mode, and is the -th learnable mode embedding.
@@ 3.6 Training Objective
To train BehaviorGPT, we apply the negative log-likelihood loss to the factorized distribution of as formulated previously:
Note that each ground-truth trajectory patch is transformed into the viewpoint of its previous patch. During training, we utilize teacher forcing to parallelize the modeling of next-patch prediction and ease the learning difficulty, but we do not use the ground-truth agent states when updating the RNN’s hidden states, intending to train the model to recover from its mistakes made in next-state prediction.
@@ 4 Experiments
This section first introduces the dataset and the evaluation metrics used in our experiments, followed by presenting the implementation details and the rollout results obtained by BehaviorGPT on the Waymo Open Sim Agents Benchmark [32]. Finally, we conduct ablation studies to further compare and analyze the performance of BehaviorGPT under various settings.
@@ 4.1 Dataset and Metrics
Our experiments are conducted on the Waymo Open Motion Dataset (WOMD) [15]. The dataset comprises 486,995/44,097/44,920 training/validation/testing scenarios. Each scenario includes -step observations sampled at Hz, totaling seconds. Given -step initial states of the scenarios, we simulate up to agents and generate simulation steps per agent at -second intervals in an autoregressive and reactive manner. Each agent requires simulations comprising x/y/z centroid coordinates and a heading value. The results on the test set are obtained by utilizing the full training set, while the performance on the validation set is based on of training data unless specified.
@@
We use various metrics for evaluation. The minADE measures the minimum average displacement error over multiple simulated trajectories, assessing trajectory accuracy. REALISM is the meta-metric that expects the simulations to match the real-world distribution. LINEAR SPEED and LINEAR ACCEL evaluate the realism regarding speed and acceleration. Similarly, ANG SPEED and ANG ACCEL measure the realism of angular speed and acceleration. DIST TO OBJ considers the distances to objects, while COLLISION and TTC assess the simulation performance in terms of collision and time to collision. Finally, DIST TO ROAD EDGE and OFFROAD focus on map compliance.
4.2 Implementation Details
The optimal patch size we experimented with is , corresponding to second. All hidden sizes are set to . Each attention layer has attention heads with dimensions per head. To save training resources, we limit the maximum number of agents per scenario to and restrict the maximum number of neighbors in kNN attention layers to . The prediction head produces modes per agent and time step. We train the models for epochs on NVIDIA RTX 4090 GPUs with a batch size of , utilizing the AdamW optimizer [31]. The weight decay rate and dropout rate are both set to . The learning rate is initially set to and decayed to following a cosine annealing schedule [30]. Our results in the 2024 WOSAC are obtained using a single model with decoding blocks and a total of 3M parameters. To produce replicas of rollouts, we randomly sample behavior modes from agents’ next-patch distributions until completing the -second multi-agent trajectories, and we repeat this process with different random seeds. The final results on the leaderboard are based on a replan rate of Hz, while the ablation studies are based on a -Hz replan rate unless specified.
4.3 Quantitative Results
@@
| Model | #Param | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
| Linear Extrapolation [32] | - | 7.5148 | 0.3985 | 0.0434 | 0.1661 | 0.2522 | 0.4393 | 0.2154 | 0.3905 | 0.7555 | 0.4801 | 0.4426 |
| TrafficBotsV1.5 [55] | 10M | 1.8825 | 0.6988 | 0.3361 | 0.3497 | 0.4512 | 0.5844 | 0.3596 | 0.8083 | 0.8209 | 0.6423 | 0.9134 |
| VBD [23] | 12M | 1.4743 | 0.7200 | 0.3591 | 0.3664 | 0.4197 | 0.5222 | 0.3683 | 0.9341 | 0.8153 | 0.6508 | 0.8788 |
| MVTE [49] | >65M | 1.6770 | 0.7302 | 0.3506 | 0.3531 | 0.4974 | 0.6000 | 0.3743 | 0.9049 | 0.8310 | 0.6655 | 0.9071 |
| GUMP [22] | 523M | 1.6041 | 0.7431 | 0.3567 | 0.4111 | 0.5089 | 0.6353 | 0.3707 | 0.9403 | 0.8276 | 0.6686 | 0.9028 |
| BehaviorGPT (Ours) | 3M | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
We report the test set results in Table 1. Notably, BehaviorGPT achieves the lowest minADE and the best REALISM, underscoring the model’s ability to match the real-world distribution. Its excellent performance on COLLISION and OFFROAD also indicates that the model has successfully captured the agent-agent and agent-map interactions in driving scenarios. Besides the benchmarking results, we also compare the number of model parameters in BehaviorGPT and other baselines. Table 1 demonstrates that BehaviorGPT, with only 3M parameters, achieves more realistic simulation than significantly larger models like MVTE [49] and GUMP [22], which demonstrates the parameter efficiency of our approach. Without employing tricks like data augmentation, model ensemble, or post-processing steps, BehaviorGPT won first place in the 2024 Waymo Open Sim Agents Challenge.
4.4 Qualitative Results
Figure 4 visualizes some qualitative results of the rollouts produced by our model. In this scenario, BehaviorGPT can generate multiple plausible futures given the same initial states of agents, which demonstrates its capability of simulating diverse yet realistic agent behavior. However, we also note that autoregressive models still suffer from accumulated errors in some cases. As shown in Figure 5, the vehicle in orange gradually goes out of the road as time goes by, which indicates the inherent limitations of autoregressive generation.
4.5 Ablation Studies
We conduct some ablation studies to gain a more in-depth understanding of our approach.
@@ 4.5 Ablation Studies
Impact of patch size. Table 2 presents the results of BehaviorGPT with varying patch sizes. According to the results, it is evident that using a patch size of , i.e., training and predicting with -Hz tokens, significantly outperforms the baseline without patching. Moreover, increasing the patch size to further enhances the overall performance. These results demonstrate the benefits of incorporating the NP3 into agent simulation. However, changing the patch size also leads to a variation in replan frequency, which also has an influence on simulation. Next, we investigate the impact of replan frequency on the test set using the model submitted to the 2024 WOSAC.
@@
| Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
| 1 | 10 Hz | 2.3752 | 0.6783 | 0.2559 | 0.2088 | 0.4022 | 0.5094 | 0.3201 | 0.9002 | 0.8015 | 0.6149 | 0.8432 |
| 5 | 2 Hz | 1.5599 | 0.7273 | 0.3543 | 0.3218 | 0.4623 | 0.5435 | 0.3768 | 0.9181 | 0.8339 | 0.6564 | 0.9077 |
| 10 | 1 Hz | 1.5203 | 0.7335 | 0.3517 | 0.3023 | 0.4734 | 0.5432 | 0.3797 | 0.9358 | 0.8329 | 0.6645 | 0.9132 |
Impact of replan frequency. During inference, we vary the replan frequency of the model with a patch size of by discarding a portion of the predicted states at each simulation step. As shown in Table 3, increasing the replan frequency from Hz to Hz can even improve the overall performance, which may benefit from the enhanced reactivity. This phenomenon demonstrates that the performance gain is not merely due to the lower replan frequency, as the model with a patch size of beats that with a patch size of even harder if using the same replan frequency of Hz. However, using an overly high replan frequency harms the performance, as indicated by the third row of Table 3. Overall, we conclude that using a larger patch indeed helps long-term reasoning, but a moderate replan frequency is important for temporal stability, which may be neglected by prior works.
@@
| Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
| 10 | 1 Hz | 1.5405 | 0.7414 | 0.3553 | 0.3153 | 0.4695 | 0.5303 | 0.3772 | 0.9520 | 0.8285 | 0.6664 | 0.9308 |
| 10 | 2 Hz | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
| 10 | 5 Hz | 1.5693 | 0.7342 | 0.3430 | 0.3472 | 0.4663 | 0.5673 | 0.3722 | 0.9429 | 0.8253 | 0.6534 | 0.9089 |
| Agent-Agent Self-Attention | minADE () | REALISM () | DIST TO OBJ () | COLLISION () | TTC () |
| 2.1489 | 0.6659 | 0.3539 | 0.6987 | 0.8070 | |
| ✓ | 1.6247 | 0.7349 | 0.3783 | 0.9409 | 0.8320 |
Impact of multi-agent interaction modeling. We remove all agent-agent self-attention layers in the first row of Table 4 to show that modeling the interactions among agents can boost minADE and REALISM. In particular, the realism in terms of collision is improved by when employing agent-agent self-attention.
| Train Data | #Param | minADE () | REALISM () |
| 20% | 5M | 1.4881 | 0.7396 |
| 50% | 5M | 1.4060 | 0.7427 |
| 100% | 5M | 1.3804 | 0.7438 |
| Model Depth | #Param | minADE () | REALISM () |
| 2 | 3M | 1.6247 | 0.7349 |
| 3 | 4M | 1.5381 | 0.7387 |
| 4 | 5M | 1.4881 | 0.7396 |
| Model Width | #Param | minADE () | REALISM () |
| 64 | 800K | 1.9637 | 0.7251 |
| 128 | 3M | 1.6247 | 0.7349 |
| 192 | 7M | 1.4993 | 0.7382 |
| Training | Inference | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
| 9.1 sec | 9.1 sec | 1.6247 | 0.7349 | 0.3546 | 0.3105 | 0.4689 | 0.5363 | 0.3783 | 0.9409 | 0.8320 | 0.6605 | 0.9163 |
| 5.0 sec | 9.1 sec | 1.6294 | 0.7333 | 0.3565 | 0.3471 | 0.4613 | 0.5293 | 0.3813 | 0.9375 | 0.8273 | 0.6585 | 0.9100 |
Scaling with data. We train our models with different proportions of training data. All the models have decoding blocks and a hidden size of , totaling 5M parameters. As shown in Table 4.5, BehaviorGPT is able to achieve remarkable performance with merely of training data, which is attributed to the high data efficiency of our approach. Increasing the proportion of training data from to further improves the performance on minADE and REALISM, and training on of the data continues to gain enhancement. Judging from the trend in Table 4.5, we believe that feeding more data for model training will continuously achieve better simulation performance.
Scaling with model size. We investigate the effects of scaling up the model size based on some preliminary experiments with of training data. In Table 4.5, we vary the number of decoding blocks while fixing the hidden size as . On the other hand, we fix the number of decoding blocks as and vary the hidden size, as depicted in Table 4.5. Based on the experimental results, we can summarize that enlarging the model consistently leads to more realistic simulation, which showcases the potential of BehaviorGPT for scaling up.
Extrapolation ability. We tried training a model on -second sequences and generating -second sequences during inference. The results in Table 8 show that this model achieves similar performance compared with the baseline trained with -second sequences, demonstrating our approach’s extrapolation ability to generate longer sequences.
@@ 5 Conclusion
This work introduced BehaviorGPT, a fully autoregressive architecture designed to enhance smart agent simulation for autonomous driving. By applying homogeneous Transformer blocks to entire trajectory snippets and utilizing relative spacetime representations, BehaviorGPT simplifies the modeling process and maximizes data utilization. To enable high-level understanding and long-range interaction reasoning in space and time, we developed the Next-Patch Prediction Paradigm, which tasks models with generating trajectory patches instead of single-step states. Experimental results on the Waymo Open Sim Agents Challenge demonstrate that BehaviorGPT achieves outstanding performance with merely 3M model parameters, highlighting its potential to further improve the realism of agent simulation with more data and computation.
Limitations. First, BehaviorGPT is currently inferior in kinematics-related performance, which can be enhanced by incorporating a kinematic model, e.g., the bicycle model. Second, the current version of BehaviorGPT does not support controlling agent behavior with specific prompts such as language and goal points. However, achieving controllable generation should be trivial given a powerful base model. Finally, we have not verified whether BehaviorGPT will facilitate the development of motion planning, which we leave as future work.
Acknowledgement
This project is supported by a grant from Hong Kong Research Grant Council under GRF project 11216323 and CRF C1042-23G.